A particle moves in a straight line with its position given by the following equation:x(t)=t^4-4t^3+2t^2+3t+6




a.Find its position after 1 second



b.Find its velocity after 2 seconds

1
Expert's answer
2021-12-15T16:21:52-0500

Given:

"x(t)=t^4-4t^3+2t^2+3t+6"


(a) the position at the time "t=1\:\rm s"

"x(1)=1^4-4*1^3+2*1^2+3*1+6=8\:\rm m"

(b) the position velocity of the particle

"v(t)=x'(t)=(t^4-4t^3+2t^2+3t+6)'\\
=4t^3-12t^2+4t+3"


at the time "t=2\:\rm s"

"v(2)=4*2^3-12*2^2+4*2+3=-5\:\rm m/s"